OpenStudy (anonymous):

I need help (AGAIN) lol Which of the following are identities? there can be more than 1 correct answers... A) cos^2 3x-sin^2 3x=cos6x B) (sinx+cosx)^2=1+sin2x C) sinx+sin5x=tan3x D) sin^6 x+cos^6 x=4-3sin^2 x

OpenStudy (anonymous):

You need to use your other identities (here is a table of them.. http://www.sosmath.com/trig/Trig5/trig5/trig5.html ) and work through each answer to find which one equals the others. Trig identities are just big puzzles; once you move the right it all works out.

OpenStudy (anonymous):

alright. thank you

OpenStudy (lgbasallote):

\[\cos^2 a - \sin^2 a = \cos (2a)\]

OpenStudy (anonymous):

(sinx+cosx)^2=sin^2x+2sinxcosx+cos^2x=1+sin2x

OpenStudy (lgbasallote):

idk what the last two mean

OpenStudy (anonymous):

mean nothing

OpenStudy (anonymous):

so you are saying the first two are correct?

OpenStudy (anonymous):

yep

OpenStudy (lgbasallote):

basically yeah

OpenStudy (anonymous):

nice. thank you